EXPERIMENTAL MAPPING OF THE ABSOLUTE VALUE OF THE ELECTRONIC TRANSITION DIPOLE MOMENT FUNCTION μ_{e}(R) OF THE ^{7}Li_{2} A^{1}Σ_{u}^{+} − X^{1}Σ_{g}^{+} SYSTEM

\underline{P. QI}, O. SALIHOGLU, E. AHMED, S. KOTOCHIGOVA, Physics Department, Temple University, Philadelphia, PA 19122; J. HUENNEKENS, Physics Department, Lehigh University, Bethlehem, PA 18015; A. M. LYYRA, Physics Department, Temple University, Philadelphia, PA 19122.

Continuous wave (CW) triple resonance laser spectroscopy^{a} was used to map the absolute value of the electronic transition dipole moment function μ_{e}(R) of the ^{7}Li_{2} A-X system. In this work, the transition dipole moment matrix elements of specific rovibronic molecular transitions were first determined by fitting the observed Autler-Townes splitting spectra. Then by employing the R-centroid approximation, or a multi-variable fit involving higher moments of R, μ_{e}(R) has been determined within a relatively large range of internuclear distance R. Finally, this electronic transition dipole moment function is compared with the ab initio calculations with very good agreement.

aE. Ahmed et al. J. Chem. Phys. 124, 084308 (2006)